The lifetime of a machine is continuous on the interval (0 ,40) with probability density function f,where f(t) is proportional to (t+10)^-2 and t is the lifetime in years.Calculate the probability that the lifetime of the machine part is less than 10 years
1
Expert's answer
2020-05-11T13:43:52-0400

"\text {Let a is a constant, then}\\
f(t)=a(t+10)^{-2}, 0<t<40\\
a\int _0^{40}\left(t+10\right)^{-2}dt=1\\
a\left[-(t+10\right)^{-1}]^{40}_{0}=1\\
-a(50^{-1}-10^{-1})=1\\
-a(\frac{-2}{25})=1\\
a=\frac{25}{2}\\
\therefore f(t)=f(t)=\frac{25}{2}(t+10)^{-2}, 0<t<40\\
P(t<10)=\frac{25}{2}\int _0^{10}\left(t+10\right)^{-2}dt\\
=\frac{25}{2}\left[-(t+10\right)^{-1}]^{10}_{0}\\
=\frac{-25}{2}(20^{-1}-10^{-1})\\
=\frac{-25}{2}\times\frac{-1}{20}\\
=\frac{5}{8}"